### v1

# Efficient Diffusion Models: A Survey

Abstract
Diffusion models have emerged as powerful generative models capable of producing high-quality contents such as images, videos, and audio, demonstrating their potential to revolutionize digital content creation. However, these capabilities come at the cost of their significant computational resources and lengthy generation time, underscoring the critical need to develop efficient techniques for practical deployment. In this survey, we provide a systematic and comprehensive review of research on efficient diffusion models. We organize the literature in a taxonomy consisting of three main categories, covering distinct yet interconnected efficient diffusion model topics from algorithm-level, system-level, and framework perspective, respectively. We have also created a GitHub repository where we organize the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-Diffusion-Model-Survey. We hope our survey can serve as a valuable resource to help researchers and practitioners gain a systematic understanding of efficient diffusion model research and inspire them to contribute to this important and exciting field.
1 Introduction
Diffusion models kickstart a new era in the field of artificial intelligence generative content (AIGC), garnering unprecedented attention (Yang et al., 2023b; Croitoru et al., 2023b). Especially in the context of image synthesis tasks, diffusion models have demonstrated impressive and diverse generative capabilities. The powerful cross-modal capabilities of diffusion models have also further fueled the vigorous development of downstream tasks (Chen et al., 2023b). Despite the increasing maturity of diffusion model variants after numerous iterations (Zhang et al., 2023d; Xu et al., 2023), generating high-resolution complex natural scenes remains both time-consuming and computationally intensive, whether the initial pixel-level approach (Ho et al., 2020) or the latent space variant (Rombach et al., 2022). Therefore, in order to optimize user-level deployment of diffusion models, researchers have never ceased their pursuit of efficient diffusion models.
Despite the growing popularity of diffusion models in recent years, one of the significant issues with diffusion model is that its multi-step denoising procedure requires numerous timesteps to reconstruct a high-quality sample from noise. This multi-step process is not only time-consuming but also computationally intensive, resulting in a heavy workload. Therefore, improving the efficiency of diffusion models is crucial. In recent years, various studies have been presented to address this problem, such as controlling the noise added during training (Hang & Gu, 2024; Chen et al., 2023a) and selecting appropriate sampling timesteps (Watson et al., 2021; Sabour et al., 2024), among others.
While there are numerous comprehensive surveys on diffusion models (Yang et al., 2023b; Chen et al., 2024; Croitoru et al., 2023a; Cao et al., 2024) and those focused on specific fields and tasks (Ulhaq et al., 2022; Lin et al., 2024c; Kazerouni et al., 2023; Lin et al., 2024b; Peng et al., 2024b; Daras et al., 2024), discussions on the efficiency of diffusion models are notably scarce. The only existing survey addressing efficiency (Ma et al., 2024c) serves as an initial exploration in this area. In our work, we provide a more comprehensive and detailed taxonomy of efficient techniques, covering a broader and more recent collection of research papers.
The overarching goal of this survey is to provide a holistic view of the technological advances in efficient diffusion models from algorithm-level, system-level, application, and framework perspectives, as illustrated in Figure 1. These four categories cover distinct yet interconnected research topics, collectively providing a systematic and comprehensive review of efficient diffusion models research. Specifically,
  • Algorithm-Level Methods: Algorithm-level methods are critical for improving the computational efficiency and scalability of diffusion models, as their training and inference processes are often resource-intensive. In §3, we survey efficient techniques that cover research directions related to efficient training, efficient fine-tuning, efficient sampling, and model compression.
  • System-Level Methods: System-level methods aim to optimize the infrastructure and computational resources required for training and deploying diffusion models. In §4, we survey efficient techniques that cover research directions related to optimized hardware-software co-design, parallel computing, and caching techniques.
  • Applications: Applications of diffusion models span various domains, where efficiency directly impacts their practical usability. Tailored techniques are required to optimize the performance of diffusion models for these specific tasks without sacrificing quality. In §2.2, we survey efficient techniques that focus on image generation, video generation, text generation, audio generation, and 3D generation.
  • Frameworks: The advent of diffusion models necessitates the development of specialized frameworks to efficiently handle their training, fine-tuning, inference, and serving. While mainstream AI frameworks such as TensorFlow and PyTorch provide the foundations, they lack built-in support for specific optimizations and features crucial for diffusion models. In §5, we survey existing frameworks specifically designed for efficient diffusion models, covering their unique features, underlying libraries, and specializations.
In addition to the survey, we have established a GitHub repository where we compile the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-Diffusion-Model-Survey. We will actively maintain it and incorporate new research as it emerges.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=12em, font=, where level=2text width=18em, font=, where level=3text width=16em, font=, where level=4text width=18em, font=, where level=5text width=18em, font=, [ Efficient Diffusion, ver [ Application (§2.2) , fill=green!10 [ Image Generation (§2.2.1), fill=green!10 ] [ Video Generation (§2.2.2), fill=green!10 ] [ Text Generation (§2.2.3), fill=green!10 ] [ Audio Generation (§2.2.4), fill=green!10 ] [ 3D Generation (§2.2.5), fill=green!10 ] ] [ Algorithm (§3), fill=blue!10 [ Efficient Training (§3.1), fill=blue!10 [ Noise Schedule, fill=blue!10 ] [ Score Matching, fill=blue!10 ] [ Data-Dependent Adaptive Priors , fill=blue!10 ] [ Rectified Flow, fill=blue!10 ] ] [ Efficient Fine-tuning (§3.2), fill=blue!10 [ LoRA, fill=blue!10 ] [ Adapter, fill=blue!10 ] [ ControlNet, fill=blue!10 ] ] [ Efficient Sampling (§3.3), fill=blue!10 [ Efficient Solver, fill=blue!10 [ SDE Solver, fill=blue!10 ] [ ODE Solver, fill=blue!10 ] ] [ Sampling Scheduling, fill=blue!10 [ Parallel Sampling, fill=blue!10 ] [ Timestep Schedule, fill=blue!10 ] ] [ Partial Sampling, fill=blue!10 [ Early Exit, fill=blue!10 ] [ Retrieval-Based Diffusion, fill=blue!10 ] ] ] [ Compression (§3.4), fill=blue!10 [ Quantization, fill=blue!10 [ Post-Training Quantization, fill=blue!10 ] [ Quantization-Aware Training, fill=blue!10 ] ] [ Pruning, fill=blue!10 ] [ Knowledge Distillation, fill=blue!10 [ Vector Field Distillation, fill=blue!10 ] [ Generator Distillation, fill=blue!10 ] ] ] ] [ System (§4), fill=orange!10 [ Hardware-Software Co-Design (§4.1) , fill=orange!10 ] [ Parallel Computing (§4.2) , fill=orange!10 ] [ Caching Technique (§4.3) , fill=orange!10 ] ] [ Frameworks (§5), fill=lime!10 [ Diffusers, DALL-E, OneDiff, LiteGen, InvokeAI, Comfy UI-Docker, Grate, DifFUSER,
Versatile Diffusion, UniDiffuser, text width=43em, fill=lime!10 ] ] ]
Figure 1: Taxonomy of efficient diffusion model literature.
2 Background and Applications
2.1 Background of Diffusion Models
To better understand the directions for improving efficient diffusion models, it is essential first to comprehend the fundamental framework of diffusion models. Denoising Diffusion Probabilistic Models (DDPMs)  (Ho et al., 2020) generate data through a process analogous to thermodynamic diffusion, consisting of two key components: a forward process and a reverse process. These processes work in concert to enable high-quality generative modeling.
The forward process in DDPM is a fixed Markov chain involving gradually adding Gaussian noise to the data until it becomes pure noise. $q(\mathbf{x}_{0})$ is denoted as the true data distribution, and assuming that $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ represents sampled data from this distribution. The forward process noted as $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$, adds Gaussian noise step by step, transforming the data from $\mathbf{x}_{0}$ to $\mathbf{x}_{T}$:
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\quad q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\beta_{t}I)$ (1)
$\beta_{t}$ is defined as the variance of the noise added at each timestep. We then convert this to $\alpha_{t}=1-\beta_{t}$. Additionally, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ is defined as the cumulative product of $\alpha_{t}$, following the formulation by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015). This cumulative product allows for modeling the transition from the original data $\mathbf{x}_{0}$ to $\mathbf{x}_{t}$ as a Gaussian distribution:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (2)
This expression describes the distribution of $\mathbf{x}_{t}$ given the initial data $\mathbf{x}_{0}$. It indicates that $\mathbf{x}_{t}$ can be expressed as a linear combination of $\mathbf{x}_{0}$ and noise, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ represents standard Gaussian noise:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ (3)
The reverse process, in contrast, aims to gradually denoise and reconstruct the original data by reversing the noise addition performed in the forward process. This reverse process is modeled as a Markov chain where each step transitions from $\mathbf{x}_{t}$ to $\mathbf{x}_{t-1}$ using a learned conditional probability distribution $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$. The overall process is expressed as:
$p_{\theta}(\mathbf{x}_{0:T}):=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}),\quad p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}):=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$ (4)
where $p(\mathbf{x}_{T})$ is the initial Gaussian distribution at the final time step $T$, and $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ represents the conditional probability distribution learned by the model to transition between states. The mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and covariance $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are parameterized functions of the state $\mathbf{x}_{t}$, the time step $t$, and the model parameters $\theta$. In the training process, the optimization objective is to minimize the negative log-likelihood using the variational bound to approximate the true data distribution:
$\mathbb{E}[-\log p_{\theta}(\mathbf{x}_{0})]\leq\mathbb{E}_{q}\left[-\log\frac{p_{\theta}(\mathbf{x}_{0:T})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right]=\mathbb{E}_{q}\left[-\log p(\mathbf{x}_{T})-\sum_{t\geq 1}\log\frac{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}\right]=:L$ (5)
This objective function decomposes the optimization problem into KL divergences for each timestep, progressively optimizing the reverse process. Expanding the KL terms and using the conditional Gaussian form evaluates the difference between the forward and reverse processes, ultimately simplifying the process into a mean squared error form:
$L_{\text{simple}}(\theta):=\mathbb{E}_{t,\mathbf{x_{0}},\bm{\epsilon}}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)\|^{2}\right]$ (6)
2.2 Applications of Diffusion Models
2.2.1 Image Generation
Image generation is the primary application domain for efficient diffusion models. Researchers have been developing various approaches to optimize both computational resources and generation quality. The efficiency improvements in this field are well exemplified by several influential works. For example, Stable Diffusion (Rombach et al., 2022) pioneered the concept of efficient image generation by operating in a compressed latent space rather than pixel space, significantly reducing memory and computational requirements while maintaining high-quality outputs. Latent Consistency Models (LCM) (Luo et al., 2023a) further pushed the boundaries by enabling high-quality image generation in just 4 steps through careful design of the consistency loss and distillation process. Progressive distillation (Salimans & Ho, 2022) demonstrated that through a student-teacher framework, diffusion models could achieve comparable quality to 50-step sampling using only 2-8 inference steps. ControlNet (Zhang et al., 2023d) introduced an efficient architecture for adding spatial conditioning controls to pretrained diffusion models through zero-initialized convolutions, enabling diverse control capabilities without compromising model efficiency. More recently, Efficient Diffusion (EDM) (Karras et al., 2022) presented a comprehensive framework for training and sampling diffusion models more efficiently, introducing improvements in both training stability and inference speed while maintaining state-of-the-art generation quality.
2.2.2 Video Generation
Following the rapid escalation in image generation, video generation similarly garnered widespread attention (Ho et al., 2022; Singer et al., 2022; Rombach et al., 2022; Xing et al., 2023). The heavy model size and the substantial computational costs have further intensified the focus on developing more efficient methods for video generation (Zhang et al., 2023a; Liu et al., 2023a; Xing et al., 2024; Wang et al., 2023; Lee et al., 2024b). For example, Zhang et al. (2023a) introduced AdaDiff, a lightweight framework designed to optimize a specialized policy gradient method tailored to individual text prompts. This approach facilitates the design of reward functions and enables an effective trade-off between inference time and generation quality. Specifically to the training process, Liu et al. (2023a) proposed an efficient training framework ED-T2V to freeze the pretraining model (Rombach et al., 2022) training additional temporal modules. Similarly, Xing et al. (2024) suggested using spatial and temporal adapters. In their approach, the original T2I model remains frozen during training, and only the newly added adapter modules are updated. Unlike the works above, Wang et al. (2023) presented VideoLCM, incorporating consistency distillation in the latent space. VideoLCM efficiently distills knowledge from a pretraining model, maintaining fidelity and temporal coherence while improving inference speed.  Lee et al. (2024b) introduces a grid diffusion model by representing a video as a grid of images. It employs key grid image generation and autoregressive grid interpolation to maintain temporal consistency.
2.2.3 Text Generation
Efficient diffusion models offer a fresh perspective in text generation through their stochastic and iterative processes. However, they encounter several challenges when applied to discrete data types such as text. For instance, the common use of Gaussian noise is less effective for discrete corruption, and the objectives designed for continuous spaces become unstable in the text diffusion process, particularly at higher dimensions. With these challenges, Chen et al. (2023a) proposed a diffusion model called Masked-Diffuse LM. In the diffusion process, a cross-entropy loss function at each diffusion step is utilized to efficiently bridge the gap between the continuous representations in the model and the discrete textual outputs. SeqDiffuSeq (Yuan et al., 2024) incorporates an encoder-decoder Transformer architecture, achieving efficient text generation through adaptive noise schedule and self-conditioning (Chen et al., 2022a) techniques. Using the same encoder-decoder architecture, Lovelace et al. (2024) presents a methodology where text is encoded into a continuous latent space. Subsequently, continuous diffusion models are employed for sampling within this space.
2.2.4 Audio Generation
In the field of audio generation, the application of diffusion models presents several unique challenges. First, audio data exhibits strong temporal continuity, especially in high-resolution audio generation tasks, where the model must accurately reconstruct both time-domain and frequency-domain information. Compared to images or text, even subtle distortions or noise in audio are easily perceptible by humans, directly affecting the listening experience, particularly in speech and music generation tasks. Ensuring high fidelity and maintaining the consistency of details in the generated audio is therefore crucial. Moreover, many audio generation tasks require low-latency feedback, especially in applications like speech synthesis and real-time dialogue, which makes acceleration of diffusion models essential. The multi-dimensional nature of audio data, such as time-domain, frequency-domain, stereo, and spatial audio, further complicates the generation process, requiring the model to handle these dimensions while maintaining consistency during the accelerated generation. To address these challenges, researchers have proposed various methods to accelerate diffusion models in audio generation. Some works focus on reducing the number of diffusion steps to speed up the generation process, such as Chen et al. (2020) in WaveGrad and Kong et al. (2020) in DiffWave, which optimize the network structure to reduce generation time while maintaining high audio quality. Further optimization comes from the FastDPM framework (Kong & Ping, 2021), which generalizes discrete diffusion steps to continuous ones, using a bijective mapping between noise levels to accelerate sampling without compromising quality. FastDPM’s flexibility allows it to adapt to different domains, and in the case of audio synthesis, where stochasticity plays a crucial role, it demonstrates superior performance in high-stochasticity tasks like speech generation. Through these approaches, diffusion models not only accelerate the generation
Table 1: Representative applications of diffusion models.
Task Datasets Metrics Articles
Image
Generation
ImageNet, CIFAR, MetFace,
CelebA HQ, MS COCO, UCI,
FFHQ, DiffusionDB, AFHQ,
LSUN, SYSTEM-X, LAION
FID, sFID, IS, NLL, MSE,
CLIP Score, PSNR, LPIPS,
MACs, CS, PickScore,
SA, Score Matching Loss
Liu et al. (2022b),
Liu et al. (2023b),
Yan et al. (2024),
Lee et al. (2024a),
Zhu et al. (2024), etc.
Video
Generation
MSR-VTT, InternVid,
WebVid-10M, LAION,
UCF-101, CGCaption
FID, IS, FVD, IQS, NIQE,
CLIPSIM, B-FVD-16
Zhang et al. (2023a),
Liu et al. (2023a),
Xing et al. (2024),
Wang et al. (2023),
Lee et al. (2024b), etc.
Audio
Generation
SC09, LJSpeech,
Speech Commands
MOS, FID, IS,
mIS, AM Score
Chen et al. (2020),
Kong et al. (2020),
Kong & Ping (2021), etc.
Text
Generation
XSUM, Semantic Content,
CCD, IWSLT14, WMT14,
ROCStories, E2E, QQP,
Wiki-Auto, Quasar-T,
AG News Topic
Rouge, Semantic Acc, Mem,
BLEU, Div, BERTScore,
SacreBLEU, MAUVE Score,
Content Fluency, POS
Chen et al. (2023a),
Yuan et al. (2024),
Chen et al. (2022a),
Lovelace et al. (2024), etc.
3D
Generation
BraTS2020, ShapeNet,
Objaverse, Cap3D, LLFF,
HumanML3D, AMASS,
KIT, HumanAct12, IBRNet,
Instruction-NeRF2NeRF
Dice, HD95, CD, EMD,
1-NNA, COV, CLIP,
Aesthetic, Similarity,
R-Precision, FID, DIV,
MM-Dist, ACC, Diversity,
MModality
Bieder et al. (2023),
Mo (2024),
Li et al. (2024c),
Park et al. (2023),
Yu et al. (2024), etc.
Table 1: (Continued)
process but also reduce computational costs while ensuring that audio quality remains high, meeting the demands of real-time audio generation applications.
2.2.5 3D Generation
As a technique closely aligned with real-world representation, 3D generation holds substantial promise across various sectors, including medical imaging, motion capture, asset production, and scene reconstruction, etc. However, when compared to 2D image generation, distinctive high-resolution elements such as volumetric data or point clouds present unique challenges, significantly escalating computational demands. Several efficient methodologies (Bieder et al., 2023; Zhou et al., ; Tang et al., 2023; Park et al., 2023) have been proposed, particularly concentrating on enhancing the sampling process and optimizing the architectural framework, which further handles the computational complexity inherent. One of the most prevalent approaches involves designing more efficient sampling schedules (Bieder et al., 2023; Li et al., 2024c; Yu et al., 2024; Zhou et al., ). By utilizing larger sampling step sizes, modifying the sampling strategy between 2D and 3D, or incorporating multi-view parallelism, these techniques address the key bottlenecks in the sampling process, thereby improving sampling efficiency. Moreover, the incorporation of novel architectures, such as state-space models and lightweight feature extractors (Mo, 2024; Tang et al., 2023), alleviates the computational burden of processing 3D data, significantly enhancing model efficiency.
3 Algorithm-Level Efficiency Optimization
3.1 Efficient Training
Efficient training techniques focus on reducing the costs of the DM pre-training process in terms of compute resources, training time, memory and energy consumption. As summarized in Figure 2, enhancing the efficiency of pre-training can be achieved through different and complementary techniques, including noise schedule, score matching, data-dependent adaptive priors, and rectified flow.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=18em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ Efficient Training, ver [ Noise Schedule, fill=blue!10 [ Fixed Noise Schedule, fill=blue!10 [ DDPM (Ho et al., 2020), IDDPM (Nichol & Dhariwal, 2021), Laplace (Hang & Gu, 2024),
Masked-Diffusion LM (Chen et al., 2023a), leaf, text width=45em ] ] [ Adaptive Noise Schedule, fill=blue!10 [ DDIM (Song et al., 2020a), ResShift (Yue et al., 2024), Immiscible Diffusion (Li et al., 2024d),
DiffuSeq (Gong et al., 2022), SeqDiffuSeq (Yuan et al., 2024), leaf, text width=45em ] ] ] [ Score Matching, fill=blue!10 [ Score Matching (Hyvärinen & Dayan, 2005), Sliced Score Matching (Song et al., 2020b), NCSN (Song & Ermon, 2019),
Likelihood Weighting (Song et al., 2021), CLD (Dockhorn et al., 2021), leaf, text width=61.7em ] ] [ Data-Dependent Adaptive Priors, fill=blue!10 [ PriorGrad (Lee et al., 2021), Digress (Vignac et al., 2022), DecompDiff (Guan et al., 2023), Leapfrog DM (Mao et al., 2023), leaf, text width=61.7em ] ] [ Rectified Flow, fill=blue!10 [ Rectified Flow (Liu et al., 2022b; Liu, 2022), InstaFlow (Liu et al., 2023b), PeRFlow (Yan et al., 2024),
2-Rectified Flows++ (Lee et al., 2024a), SlimFlow (Zhu et al., 2024), leaf, text width=61.7em ] ] ]
Figure 2: Summary of efficient training techniques for diffusion models.
3.1.1 Noise Schedule
Noise schedule is a crucial component of diffusion models, governing how noise is added at each step of the forward process and how it is removed during the reverse process. Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) first proposed a linear noise schedule that gradually decreases the variance of the noise added in the forward process in  Eq.(1). However, despite DDPM demonstrating significant potential in diffusion model tasks, this linear schedule requires calculating complex noise terms and iterating through numerous timesteps, posing challenges to noise injection efficiency and highlighting the need for an efficiency-enhanced noise schedule design. Current studies still face prolonged iterative processes, significantly hindering diffusion speed, while several studies have concentrated on designing the highlight of noise schedules. As shown in Figure 3, efficient noise schedules can be classified into fixed noise schedule and adaptive noise schedule.
Figure 3: Illustration of two categories of noise schedules.
Fixed Noise Schedule. Fixed noise schedules refer to a technique where noise is systematically added to a model’s training process at predefined intervals or according to specific levels. DDPM (Ho et al., 2020) uses a linear noise schedule, serving as a classic example of fixed noise schedules, with noise variance changing deterministically over time. Based on this, the Improved Denoising Diffusion Probabilistic Model (IDDPM) (Nichol & Dhariwal, 2021) introduces significant innovations in the noise schedule, replacing the linear noise schedule of the original DDPM with a cosine noise schedule, which could be written as
$\beta_{t}=1-\frac{\cos\left(\frac{t/T+s}{1+s}\cdot\frac{\pi}{2}\right)}{\cos\left(\frac{s}{1+s}\cdot\frac{\pi}{2}\right)}$ (7)
where $t$ is the current timestep, $T$ is the total number of timesteps, and $s$ is a small positive constant typically used to smooth the initial noise addition. The cosine noise schedule helps the model achieve high-quality sample generation by optimizing the noise distribution throughout the training process, allowing the model to generate samples with a reduced number of sampling steps while preserving essential structural information at each stage. However, the cosine noise schedule allocates computational resources evenly across the entire range of noise intensities. This means that both high and low noise samples consume significant resources, even though they may not be the most critical for training the model.
To address the inefficiency in resource allocation,  (Hang & Gu, 2024) proposed a fixed noise schedule, called Laplace. This schedule aims to further optimize the training of the diffusion model by increasing the sampling frequency around critical regions. This approach ensures that the model receives more training samples in these crucial regions, thereby improving overall training efficiency and sample quality. By redesigning the fixed noise schedule, more computational resources are concentrated on medium noise intensities. These medium noise samples are more important for model learning, as they effectively train the model to understand the data structure and remove noise. Laplace effectively balances noise addition across all time steps, ensuring a more robust and consistent training process, and ultimately resulting in higher-quality generated images.
Lastly, existing diffusion models have employed a uniform addition of Gaussian noise to each word to facilitate the diffusion process for text generation. However, this approach has been shown ta failure to leverage the linguistic features of the text in question, while simultaneously increasing the computational burden. Furthermore,  (Chen et al., 2023a) addresses this issue by gradually introducing noise through a soft-masking noise strategy. This model determines the masking order based on word importance, which is measured by term frequency and information entropy. The model employs a square-root noise schedule(Li et al., 2022) to incrementally increase the noise level, thereby stabilizing the training process. Consequently, the model gradually adds noise to the initial latent variable $X_{0}$, resulting in a series of noisy latent variables $X_{1:T}$.
Adaptive Noise Schedule. Different from the fixed noise schedules, adaptive noise schedules include methods that dynamically adjust the noise schedule based on the state of the model or data. The key idea is to adapt the noise schedule in response to specific conditions or states during the diffusion process.  Song et al. (2020a) developed Denoising Diffusion Implicit Models (DDIM), which improve the noise schedule in DDPM by introducing non-Markovian forward processes. as follows:
$x_{t-1}=\sqrt{\alpha_{t-1}}\left(x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t})\right)+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(x_{t})+\sigma_{t}\epsilon_{t}$ (8)
where $\alpha_{t}$ is a decreasing schedule controlling how noise is added over time. They propose a dynamic method for adjusting the noise level, $\sigma_{t}$, at each step. The noise level is calculated as a function of both the current state, $x_{t}$, and the initial state, $x_{0}$, leveraging the entire trajectory rather than treating each step independently. This dynamic adjustment fully utilizes the information from the initial and current states, allowing for more accurate control of the diffusion process. As a result, the quality of the generated samples is enhanced, and the number of sampling steps required is reduced.
Inspired by DDIM,  Yue et al. (2024) proposed ResShift, introducing a novel noise schedule that constructs a much shorter Markov chain. The key idea is to shift the residual between the high-resolution (HR) and low-resolution (LR) images instead of adding Gaussian noise. This approach allows the model to start from an LR image and iteratively refine it to produce the HR image. The noise schedule also involves a hyperparameter $\kappa$, which controls the overall noise intensity during the transition. The mathematical formulation of this noise schedule is as follows:
$\sqrt{\eta_{t}}=\sqrt{\eta_{1}}\times b_{0}^{\beta_{t}},\quad t=2,\ldots,T-1$ (9)
where
$\beta_{t}=\left(\frac{t-1}{T-1}\right)^{p}\times(T-1),\quad b_{0}=\exp\left(\frac{1}{2(T-1)}\log\frac{\eta_{T}}{\eta_{1}}\right)$ (10)
where $T$ is the total number of timesteps, $t$ is the current timestep, $p$ is a hyperparameter controlling the growth rate of $\sqrt{\eta_{t}}$, and $\eta_{1}$ and $\eta_{T}$ represent the initial and final noise levels, respectively. This formula enables a non-uniform geometric progression of noise levels. Thus ResShift can achieve high-quality SR results with only 15 sampling steps. Conventional methodologies tend to diffuse each image across the entirety of the noise space, thereby resulting in a composite of all images at each point within the noise layer. In light of the aforementioned,  Li et al. (2024d) proposed Immiscible Diffusion, a novel approach inspired by the physical phenomenon of immiscibility. In contrast to conventional methodologies, Immiscible Diffusion involves reassigning noise to images in a way that minimizes the distance between image-noise pairs within a mini-batch, ensuring that each image is only diffused to nearby noise points. This method ensures that each image is matched only with nearby noise, thereby reducing the difficulty of denoising. Previous studies  Gong et al. (2022) on text generation have employed fixed noise schedules and encoder-only Transformer architectures. This approach necessitated the recalculation of the input sequence at each time step, resulting in the inefficient utilization of computational resources and the generation of outputs at a slow pace. In contrast,  Yuan et al. (2024) introduces an adaptive noise scheduling technique that enables the dynamic adjustment of the noise level at each time step and token position. In particular, this technique entails recording loss values at each timestep throughout the training phase, with a linear interpolation subsequently employed to map these losses to the corresponding noise schedule parameters.
3.1.2 Score Matching
Score matching, introduced by Hyvärinen & Dayan (2005), serves as an effective approach for estimating unnormalized models by minimizing the Fisher divergence between the score function of data distribution $\bm{s}_{d}(\mathbf{x})=\nabla_{\mathbf{x}}\log p_{d}(\mathbf{x})$ and the score function of model distribution $\bm{s}_{m}(\mathbf{x};\bm{\theta})=\nabla_{\mathbf{x}}\log p_{m}(\mathbf{x};\bm{\theta})$. This approach avoids the need to compute the intractable partition function $\bm{Z}_{\bm{\theta}}$, a common problem in Maximum Likelihood Estimation (MLE).
While DDPM directly optimizes the noise prediction in Eq.(6), score matching objectives can directly be estimated on a dataset and optimized with stochastic gradient descent, the loss function for score matching takes a different approach formulated as follows:
$L(\bm{\theta})=\frac{1}{2}\mathbb{E}_{p_{d}(\mathbf{x})}\left[\|\bm{s}_{m}(\mathbf{x};\bm{\theta})-\bm{s}_{d}(\mathbf{x})\|^{2}\right]$ (11)
Since it typically does not have access to the true score function of the data $\bm{s}_{d}(\mathbf{x})$Hyvärinen & Dayan (2005) introduced integration by parts as $L(\bm{\theta})=J(\bm{\theta})+C$ to derive an alternative expression that does not require direct access to $\mathbf{x}_{d}(\mathbf{x})$:
$J(\bm{\theta})=\mathbb{E}_{p_{d}(\mathbf{x})}\left[\text{tr}(\nabla_{\mathbf{x}}\bm{s}_{m}(\mathbf{x};\bm{\theta}))+\frac{1}{2}\|\bm{s}_{m}(\mathbf{x};\bm{\theta})\|^{2}\right]$ (12)
In this expression, $\text{tr}(\cdot)$ denotes the trace of the Hessian matrix of $\bm{s}_{m}(\mathbf{x};\bm{\theta})$. The constant $C$ is independent of $\bm{\theta}$ and can be ignored for optimization purposes. The final form of the unbiased estimator used for training is:
$\hat{J}(\bm{\theta};\mathbf{x}_{1}^{N})=\frac{1}{N}\sum_{i=1}^{N}\left[\text{tr}(\nabla_{\mathbf{x}}\bm{s}_{m}(\mathbf{x}_{i};\bm{\theta}))+\frac{1}{2}\|\bm{s}_{m}(\mathbf{x}_{i};\bm{\theta})\|^{2}\right]$ (13)
Compared to DDPM’s straightforward optimization in Eq.(6), although score matching Eq.(13) avoids the computation of the partition function $\bm{Z}_{\bm{\theta}}$, it still faces computational challenges, particularly in high-dimensional data. The computation of the trace of the Hessian matrix substantially increases the complexity as the dimensionality grows. Specifically, computing the trace requires many more backward passes than the gradient, making score matching computationally expensive for high-dimensional data.
To address this issue, Song et al. (2020b) observed that one-dimensional problems are typically much easier to solve than high-dimensional ones. Inspired by the idea of the Sliced Wasserstein Distance (Rabin et al., 2012), they proposed Sliced Score Matching. The core idea of sliced score matching is to project both the score function of the model $\bm{s}_{m}(\mathbf{x};\bm{\theta})$ and the data $\bm{s}_{d}({\mathbf{x}})$ onto a random direction $\mathbf{v}$, and compare the differences along that direction. The sliced score matching objective is defined as:
$L(\bm{\theta};p_{\mathbf{v}})=\frac{1}{2}\mathbb{E}_{p_{\mathbf{v}}}\mathbb{E}_{p_{d}(\mathbf{x})}\left[\left(\mathbf{v}^{\top}\bm{s}_{m}(\mathbf{x};\bm{\theta})-\mathbf{v}^{\top}\bm{s}_{d}({\mathbf{x}})\right)^{2}\right]$ (14)
To eliminate the dependence on $\bm{s}_{d}({\mathbf{x}})$, integration is applied by parts, similar to traditional score matching, resulting in the following form:
$J(\bm{\theta};p_{\mathbf{v}})=\mathbb{E}_{p_{\mathbf{v}}}\mathbb{E}_{p_{d}(\mathbf{x})}\left[\mathbf{v}^{\top}\nabla_{\mathbf{x}}\bm{s}_{m}(\mathbf{x};\bm{\theta})\mathbf{v}+\frac{1}{2}(\mathbf{v}^{\top}\bm{s}_{m}(\mathbf{x};\bm{\theta}))^{2}\right]$ (15)
which achieves scalability by reducing the complexity of the problem by projecting high-dimensional score functions onto low-dimensional random directions, thereby avoiding the full Hessian computation.
Figure 4: Illustration of data-dependent adaptive priors for diffusion processes across different modalities, demonstrating how tailored priors improve generate on quality.
While effective for dimensionality reduction, score estimation still faces challenges in low data density regions where data samples are sparse. Building upon sliced score matching, to address the issue of inaccurate score estimation in low data density regions, Song & Ermon (2019) introduces a novel generative framework that employs Langevin dynamics to produce samples based on estimated gradients of the data distribution $p_{\text{data}}(\mathbf{x})$. They proposed Noise Conditional Score Networks (NCSN) $s_{\theta}(\mathbf{x},\sigma)$, which jointly estimate scores across multiple noise-perturbed data distributions. By conditioning on a geometric sequence of noise levels $\sigma_{3}>\sigma_{2}>\sigma_{1}$, a single network learns to estimate scores for distributions ranging from highly smoothed $p_{\sigma_{3}}(\mathbf{x})$ that fill low-density regions to concentrated $p_{\sigma_{1}}(\mathbf{x})$ that preserve the structure of the original data manifold. This unified training approach enables robust score estimation across the entire data space.
Different from sliced score matching, Song et al. (2021) introduces a new weighting scheme called likelihood weighting, which allows the weighted combination of score matching losses to actually bound the negative log-likelihood of the diffusion models. Compared with traditional score matching, it focuses on adjusting the training objective. This modification enables approximate maximum likelihood training, rather than merely minimizing the score matching loss, which improves the model’s likelihood across various datasets, stochastic processes, and architectures, which effectively combines the training efficiency of score matching with the advantages of maximum likelihood estimation. Following a similar derivation, as Song et al. (2021), Dockhorn et al. (2021) introduces Coupled Langevin Dynamics (CLD), redefining the score matching objective within the CLD framework. Unlike traditional score matching methods that inject noise directly into the data space, CLD simplifies the task by only requiring the model to learn the score of the conditional distribution $p_{t}(v_{t}\mid x_{t})$, where noise is injected into an auxiliary variable $v_{t}$ coupled with the data.
3.1.3 Data-Dependent Adaptive Priors
To accelerate the generation process in diffusion models and improve the quality of generated samples, better initialization of priors tailored to specific tasks and datasets can be utilized. This approach leverages data-dependent adaptive priors, making the generation process more efficient and aligning the generated samples more closely with true data distribution. Recent studies have explored how data-dependent adaptive priors can enhance diffusion models.
As illustrated in Figure 4, data-dependent adaptive priors can be applied across different modalities, such as speech, graph, and trajectory, to enhance the diffusion process. By aligning priors with data distributions specific to each modality, the model can generate outputs that are better tailored to the underlying structure of the data. In traditional diffusion models, the prior is typically assumed to be a standard Gaussian distribution $p(z)=\mathcal{N}(0,I)$. However, this may not always align well with the actual data distribution, potentially introducing inefficiencies. By leveraging data-dependent adaptive priors based on $X$, where $X$ represents the data, the model can achieve better initialization and faster convergence, without relying solely on a standard Gaussian assumption. Lee et al. (2021) introduce PriorGrad, which improves diffusion models for speech synthesis by using an adaptive prior derived from data statistics based on conditional information. This method increases the efficiency of the denoising process, leading to faster convergence and inference speed, while also improving perceptual quality and robustness with smaller network capacities.
Digress (Vignac et al., 2022) presents a discrete denoising diffusion model focused on graph generation. This model leverages data-dependent priors to better capture the discrete nature of graph data, significantly enhancing the quality of generated graphs, particularly for applications such as chemical molecular structures and social networks. DecompDiff (Guan et al., 2023) introduce decomposed priors, modeling different structural components of drug molecules separately to improve diffusion models. This approach enhances the accuracy of generating viable drug candidates by better capturing molecular structure information. As illustrated in Figure 5, Mao et al. (2023) propose the Leapfrog Diffusion Model for stochastic trajectory prediction, introducing a leapfrog initializer that uses adaptive priors to skip multiple denoising steps. This method significantly accelerates inference while maintaining high prediction accuracy, making it useful for real-time applications such as autonomous driving and robotic navigation.
Figure 5: The Leapfrog diffusion model (Mao et al., 2023) accelerates inference by using a leapfrog initializer to approximate the denoised distribution, replacing extended denoising sequences while preserving representation capacity.
3.1.4 Rectified Flow
Figure 6: Illustration of the rectified flow.
As illustrated in Figure 6, Rectified Flow, proposed by (Liu et al., 2022b; Liu, 2022), introduces a method for training ordinary differential equation (ODE) models by learning straight transport paths between two distributions, $\pi_{0}$ and $\pi_{1}$. The key idea is to minimize the transport cost by ensuring that the learned trajectory between these two distributions follows the most direct route, which can be computationally efficient to simulate. Unlike traditional diffusion models, which may follow roundabout paths, Rectified Flow leverages a simpler optimization problem to create a straight flow, thereby improving both training efficiency and the quality of the generated samples.
Building upon this foundation, InstaFlow (Liu et al., 2023b) apply the Rectified Flow concept to text-to-image generation, achieving a significant breakthrough. InstaFlow represents a major advancement in efficient diffusion models, which are capable of high-quality image generation in just one step. It applied Rectified Flow to large-scale datasets and complex models like Stable Diffusion, introduced a novel text-conditioned pipeline for one-step image generation, and achieved an FID score of 23.3 on MS COCO 2017-5k. InstaFlow’s success highlights the potential of Rectified Flow in dramatically reducing the computational cost of diffusion models while maintaining high output quality.
Following InstaFlow, Yan et al. (2024) proposed PeRFlow, further extending the Rectified Flow concept to create a more flexible and universally applicable acceleration method. PeRFlow divides the sampling process into multiple time windows, applying the reflow operation to each interval, creating piecewise linear flows that allow for more nuanced trajectory optimization. Through carefully designed parameterizations, PeRFlow models can inherit knowledge from pretrained diffusion models, achieving fast convergence and superior transfer ability. This approach positions PeRFlow as a universal plug-and-play accelerator compatible with various workflows based on pretrained diffusion models. While Rectified Flow showed great promise, there was still room for improvement, especially in low Number of Function Evaluations (NFE) settings. Addressing this, Lee et al. (2024a) focused on enhancing the training process of Rectified Flows. They discovered that a single iteration of the Reflow algorithm is often sufficient to learn nearly straight trajectories and introduced a U-shaped timestep distribution and LPIPS-Huber premetric to improve one-round training. These improvements led to significant enhancements in FID scores, particularly in low NFE settings, outperforming state-of-the-art distillation methods on various datasets. Most recently, Zhu et al. (2024) proposed SlimFlow, a method designed to address the joint compression of inference steps and model size within the Rectified Flow framework, introducing Annealing Reflow to address initialization mismatches between large teacher models and small student models, and developing Flow-Guided Distillation to improve performance on smaller student models.
3.2 Efficient Fine-Tuning
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10em,font=,, where level=2text width=30em,font=,, where level=3text width=30em,font=,, [ Efficient Fine-tuning [ LoRA, fill=blue!10 [ LoRA (Hu et al., 2021), LCM-LoRA (Luo et al., 2023b),
Concept Sliders (Gandikota et al., 2023), LoRA-Composer
 Yang et al. (2024), LoRA conditioning (Choi et al., 2024b), ] ] [ Adapter, fill=blue!10 [ T2I-Adapter (Mou et al., 2024), IP-Adapter (Ye et al., 2023),
ParaTAA (Tang et al., 2024), CTRL-Adapter
 (Lin et al., 2024a), SimDA (Xing et al., 2024), ] ] [ ControlNet, fill=blue!10 [ ControlNet (Zhang et al., 2023c), ControlNet++ (Li et al., 2025),
ControlNet-XS (Zavadski et al., 2023), ControlNeXt
 (Peng et al., 2024a), Uni-ControlNet (Zhao et al., 2024),
UniControl (Qin et al., 2023), ] ] ]
Figure 7: Summary of efficient fine-tuning techniques for diffusion models.
Fine-tuning pre-trained diffusion models can be computationally expensive. To address this, we categorize efficient fine-tuning techniques into LoRA, Adapters, and ControlNet, based on their mechanisms for reducing resource consumption. This classification reflects differences in how these methods optimize parameters, enable task-specific adaptation, and incorporate spatial conditioning signals, as summarized in Figure 7.
3.2.1 LoRA
Low-Rank Adaptation (LoRA) (Hu et al., 2021) is a model adaptation method that maintains frozen pre-trained model weights while enabling efficient task adaptation through the injection of low-rank decomposition matrices into each Transformer layer. The core mathematical foundation of this approach lies in its representation of the weight update mechanism: for a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA represents the weight update as:
$W=W_{0}+\Delta W,\text{ where }\Delta W=BA$ (16)
where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ are trainable low-rank matrices, and the rank $r\ll\min(d,k)$. During forward propagation, for an input $x\in\mathbb{R}^{k}$, the model computes the hidden representation $h\in\mathbb{R}^{d}$ as:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (17)
The complete process is illustrated in Figure 8. A key advantage of this design lies in its deployment efficiency, where the explicit computation and storage of $W=W_{0}+BA$ enables standard inference procedures without introducing additional latency. Originally proposed for fine-tuning Large Language Models (LLMs), LoRA has demonstrated remarkable parameter efficiency and memory reduction in model adaptation.
Figure 8: Illustration of Hu et al. (2021)’s reparameterization approach, where only parameters A and B are trained.
While predominantly utilized in LLM fine-tuning, recent research has extended its application to diffusion models, indicating its potential as a versatile adaptation technique across different deep learning architectures. LCM-LoRA (Luo et al., 2023b) proposes a universal acceleration approach for diffusion models. As shown in Figure 9, this method achieves fast sampling by adding an Acceleration vector $\tau_{LCM}$ to the Base LDM Rombach et al. (2022). This module adopts LoRA (Hu et al., 2021) to attach low-rank matrices to the original model without architectural modifications. For customized diffusion models that are fine-tuned for specific text-to-image generation tasks, the task-specific LoRA ($\tau^{\prime}$) and acceleration LoRA ($\tau_{LCM}$) can be linearly combined through Eq.(18) to achieve fast inference while maintaining generation quality. More importantly, it provides a plug-and-play solution that reduces sampling steps from dozens to around 4, while maintaining compatibility with any pre-trained text-to-image diffusion model.
$\tau^{\prime}_{LCM}=\lambda_{1}\tau^{\prime}+\lambda_{2}\tau_{LCM}$ (18)
Beyond the acceleration achieved by LCM-LoRA, Concept Sliders (Gandikota et al., 2023) extends LoRA’s application to precise control over image generation attributes. This method identifies low-rank directions in the diffusion parameter space corresponding to specific concepts through LoRA adaptation. The method freezes the original model parameters and trains a LoRA adapter to learn concept editing directions. Given an input $(x_{t},c_{t},t)$, where $x_{t}$ is the noisy image at timestep $t$.
Figure 9: Illustration of LCM-LoRA (Luo et al., 2023b).
For a target concept $c_{t}$, the model is guided through a score function to enhance certain attributes $c_{+}$ while suppressing others $c_{-}$. This training objective can be formulated as:
$\epsilon_{\theta^{*}}(x,c_{t},t)\leftarrow\epsilon_{\theta}(x,c_{t},t)+\eta(\epsilon_{\theta}(x,c_{+},t)-\epsilon_{\theta}(x,c_{-},t))$ (19)
where $\epsilon_{\theta}$ represents the denoising model’s prediction, and $\eta$ is a guidance scale. With this formulation, the method enables smooth control over concept strength through the guidance scale $\eta$ while maintaining concept independence in the learned directions. By leveraging LoRA’s parameter-efficient nature, it achieves precise attribute manipulation with minimal computational overhead.
Besides, LoRA-Composer Yang et al. (2024) advances LoRA’s application in diffusion models toward seamless multi-concept integration. While previous works focus on acceleration or single-concept control, this approach tackles the challenging task of combining multiple LoRA-customized concepts within a single image generation process. It combines multiple LoRAs in diffusion models by modifying the U-Net’s attention blocks. Specifically, it enhances both cross-attention and self-attention layers within U-Net to enable direct fusion of multiple LoRAs. Compared to traditional methods like Mix-of-Show Gu et al. (2024) that require training a fusion matrix to merge multiple LoRAs, which increases computational overhead and may degrade generation quality. It directly combines multiple lightweight LoRAs through modified attention blocks, avoiding the overhead of retraining models. While LoRA-Composer focuses on fusing multiple LoRAs for multi-concept control, Choi et al. (2024b) explores the fundamental application of LoRA in attention layers. Both these works enhance diffusion models by modifying the attention mechanism in U-Net. The latter proposes a structured conditioning approach in U-Net blocks with three key components: (1) conventional convolutional blocks using scale-and-shift conditioning for feature normalization adjustment, (2) attention blocks enhanced by LoRA adapters that condition both QKV computation and projection layers through learnable low-rank matrices, and (3) two LoRA conditioning implementations - TimeLoRA/ClassLoRA for discrete-time settings and UC-LoRA for continuous SNR settings, which utilize MLP-generated weights to combine multiple LoRA bases. Them method achieves improved performance over traditional conditioning while only increasing the parameter count by approximately 10% through efficient low-rank adaptations in the attention layers.
3.2.2 Adapter
Figure 10: Architecture of the Adapter module: demonstrating the integration of adapter layers within a transformer block to achieve efficient task adaptation by adding lightweight transformations, while keeping the core model weights frozen.
Adapters are lightweight modules designed to enable efficient task adaptation by introducing small network layers into pre-trained models, allowing task-specific feature learning while keeping the original weights frozen. As illustrated in Figure 10, adapter layers are placed within the transformer block, positioned between normalization and feed-forward layers. Each adapter module consists of a down-projection, nonlinearity, and up-projection, which generates task-specific transformations without altering the core model’s structure. This design significantly reduces memory and computational requirements, making adapters well-suited for tasks requiring lightweight parameter updates, such as text-to-image generation (T2I) and simulated domain adaptation (SimDA).
T2I-Adapter (Mou et al., 2024) is an adapter designed to enhance control in text-to-image generation models by introducing conditional features such as sketches, depth maps, and semantic segmentation maps, allowing for structural guidance in generated images. Unlike approaches that require modifying the model’s core architecture, T2I-Adapter uses lightweight modules to incorporate external condition information into the generation process without altering the pre-trained model itself. This method improves the accuracy and consistency of generated images without increasing computational costs. In implementation, T2I-Adapter employs convolutional and residual blocks to align conditional inputs with the spatial dimensions of intermediate features in the UNet model, thus capturing structural information at multiple scales. This integration allows T2I-Adapter to flexibly incorporate conditional features, such as sketches and depth maps, providing enhanced control over text-to-image generation. Such multi-adapter strategies are particularly suitable for tasks requiring high customization in image generation, enabling simultaneous input of various structural features to refine the output.
IP-Adapter (Ye et al., 2023) enhances the consistency and visual quality of text-to-image generation by incorporating image prompts. Unlike T2I-Adapter (Mou et al., 2024), which provides structural guidance through sketches or depth maps, IP-Adapter focuses on capturing visual details from an input image, making it ideal for tasks requiring high visual consistency with a reference image. This adapter processes the input image prompt into latent features, allowing the generation model to capture visual information from the target image and maintain detail alignment throughout the generation process. In its workflow, the image prompt is first mapped into the latent space and then processed through convolution and normalization modules within the adapter, enabling the model to utilize these features during inference. This setup enables the generation model to draw rich visual information from the image prompt, making IP-Adapter particularly suitable for tasks requiring high detail consistency, such as generating images with a style similar to the input image.
Figure 11: Architecture of IP-Adapter (Ye et al., 2023) using a decoupled cross-attention strategy, where only newly added modules are trained, and the pre-trained text-to-image model remains frozen.
CTRL-Adapter (Lin et al., 2024a) is designed to enhance attribute control during generation by guiding specific attributes such as emotion or object type, enabling precise customization in generated results. Unlike T2I-Adapter (Mou et al., 2024) and IP-Adapter (Ye et al., 2023), which focus on structural and detail consistency respectively, CTRL-Adapter is tailored to provide diversity control for the generation model. For example, as illustrated in Figure 11, the IP-Adapter architecture employs a decoupled cross-attention strategy, where only newly added modules are trained while the pre-trained text-to-image model remains frozen. In contrast, CTRL-Adapter can adjust the style of generated images based on specified emotions or object types, achieving controllable content generation without altering the core architecture of the model. This makes CTRL-Adapter particularly suitable for tasks requiring high customization in generation, such as emotion-driven text generation or stylized image synthesis.
SimDA (Xing et al., 2024) is an adapter suited for cross-domain generation tasks, achieving domain adaptation by utilizing simulated data within the adapter to enhance the model’s performance on previously unseen data distributions. Unlike CTRL-Adapter (Lin et al., 2024a), which primarily focuses on attribute control, SimDA is designed to improve the model’s generalization ability, allowing it to generate high-quality content even in unfamiliar data environments. SimDA is particularly useful in generation tasks that require domain transfer, such as adapting a model trained on one image dataset to perform well on another dataset. This enables the model to align with new data characteristics without compromising generation quality.
3.2.3 ControlNet
ControlNet (Zhang et al., 2023c) and its derivatives represent a significant advancement in adding spatial conditioning controls to pre-trained text-to-image diffusion models. The original ControlNet architecture (Zhang et al., 2023c), as illustrated in Figure 12, presents a novel approach to integrating various spatial conditions—such as scribbles, edge maps, open-pose skeletons, or depth maps—into the generative process while preserving the robust features of pre-trained diffusion models. The architecture employs zero convolution layers that gradually develop parameters without disrupting the pre-trained model’s stability. This design enables versatile conditioning, allowing the model to effectively leverage different types of spatial information. Through these conditioning methods, ControlNet demonstrates a remarkable ability to guide generation with fine-grained control over structure, style, and composition. Building upon this foundation, several works have proposed improvements and alternatives. ControlNet++ (Li et al., 2025) addresses the challenge of alignment between generated images and conditional controls by introducing pixel-level cycle consistency optimization. Through a pre-trained discriminative reward model and an efficient reward strategy involving single-step denoised images, it achieves significant improvements in control accuracy, with notable gains in metrics such as mIoU (11.1%), SSIM (13.4%), and RMSE (7.6%) across various conditioning types. ControlNet-XS (Zavadski et al., 2023) reimagines the control system by enhancing the communication bandwidth between the controlling network and the generation process. This redesign not only improves image quality and control fidelity but also significantly reduces the parameter count, resulting in approximately twice the speed during both inference and training while maintaining competitive performance in pixel-level guidance tasks. The field has also seen efforts to unify multiple control capabilities. UniControl (Qin et al., 2023) introduces a task-aware HyperNet approach that enables a single model to handle diverse visual conditions simultaneously. Similarly, Uni-ControlNet (Zhao et al., 2024) proposes a unified framework supporting both local controls and global controls through just two additional adapters, significantly reducing training costs and model size while maintaining high performance.
Figure 12: Illustration of ControlNet.
Most recently, ControlNeXt (Peng et al., 2024a) has pushed the boundaries of efficiency even further by introducing a streamlined architecture that minimizes computational overhead. It replaces the traditional heavy additional branches with a more concise structure and introduces Cross Normalization (CN) as an alternative to zero convolutions. This approach achieves fast and stable training convergence while reducing learnable parameters by up to 90% compared to previous methods.
3.3 Efficient Sampling
Efficient sampling in diffusion models addresses the computational challenges of the iterative denoising process while maintaining the quality of the generated samples through three main approaches. As illustrated in Figure 13, these encompass efficient SDE and ODE solvers, sampling scheduling strategies including parallel sampling and timestep optimization, and truncated sampling methods leveraging early exit and retrieval-based techniques.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=18em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ Efficient Sampling, ver [ Efficient Solver, fill=blue!10 [ SDE Solver, fill=blue!10 [ DiffFlow (Zhang & Chen, 2021), DiNoF (Zand et al., 2023),  Jolicoeur-Martineau et al. (2021),
GMS (Guo et al., 2024), BFM-Solver2 (Xue et al., 2024a), SA-Solver (Xue et al., 2024b), leaf, text width=45em ] ] [ ODE Solver, fill=blue!10 [ DDIM (Song et al., 2020a), DPM-solver (Lu et al., 2022), DEIS (Zhang & Chen, 2022),
DMCMC (Kim & Ye, 2022), i-DODE (Zheng et al., 2023), leaf, text width=45em ] ] ] [ Sampling Scheduling, fill=blue!10 [ Parallel Sampling, fill=blue!10 [ DEQ (Pokle et al., 2022), ParaDiGMS (Shih et al., 2024), ParaTAA (Tang et al., 2024),
TAA (Walker & Ni, 2011), leaf, text width=45em ] ] [ Timestep Scheduling, fill=blue!10 [ FastDPM (Kong & Ping, 2021),  Watson et al. (2021), AYS (Sabour et al., 2024), leaf, text width=45em ] ] ] [ Truncated Sampling, fill=blue!10 [ Early Exit, fill=blue!10 [ CATs (Schuster et al., 2021), ASE (Moon et al., 2024), leaf, text width=45em ] ] [ Retrieval-Guided Initialization, fill=blue!10 [ RDM (Blattmann et al., 2022), kNN-Diffusion (Sheynin et al., 2022), Re-Imagen (Chen et al., 2022b),
ReDi (Zhang et al., 2023b), leaf, text width=45em ] ] ] ]
Figure 13: Summary of efficient sampling techniques for diffusion models.
3.3.1 Efficient Solver
Given that the cost of sampling escalates proportionally with the number of discretized time steps, many researchers have concentrated on devising discretization schemes that reduce the number of time steps. A key insight emerges from reexamining the discrete forward process in the original DDPM formulation Eq.(1), as we reduce the step size between consecutive steps, the process naturally approaches a continuous transformation. Consequently, adopting learning-free methods using SDE or ODE solvers Song et al. (2020c) has been proposed.
SDE Solver. (Song et al., 2020c) firstly presents a stochastic differential equation (SDE) that smoothly transforms a complex data distribution to a known prior distribution by slowly injecting noise and a corresponding reverse-time SDE that transforms the prior distribution back into the data distribution by slowly removing the noise.The discrete noise addition steps in Eq.(1) are reformulated into a continuous process:
SDE accomplishes the transformation from data to noise in the diffusion training process through the following equation:
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t)\,dt+g(t)\,d\bar{\mathbf{w}}$ (20)
where $\bar{\mathbf{w}}$ denotes the standard Wiener process, also known as Brownian motion. $\mathbf{f}(\mathbf{x},t)$ is a vector-valued function called the drift coefficient of $\mathbf{x}(t)$, and $g(t)$ is a scalar function.
Similarly, the reverse process Eq.(4) can be generalized to a continuous-time formulation:
$d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\right]dt+g(t)d\bar{\mathbf{w}}$ (21)
$\bar{\mathbf{w}}$ is a standard Wiener process when time flows backward from $T$ to 0, $dt$ is an infinitesimal negative timestep and $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ represent the score function that we mentioned in Eq.(20). In the diffusion process, reverse-time SDE converts noise into data gradually. The complete SDE process is shown in Figure 14.
Figure 14: Overview of forward SDE process and reverse SDE process (Song et al., 2020c).
Subsequently, there are many ways to efficiently implement SDE-based solvers. (Zhang & Chen, 2021)introduces a novel generative modeling and density estimation algorithm called Diffusion Normalizing Flow (DiffFlow). Similar to the SDE of diffusion models Eq.(20), the DiffFlow model also has a forward process:
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t,\theta)dt+g(t)d\bar{\mathbf{w}}$ (22)
and a backward process:
$d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t,\theta)-g^{2}(t)\mathbf{s}(\mathbf{x},t,\theta)\right]dt+g(t)d\bar{\mathbf{w}}$ (23)
As a result of the learnable parameter $\theta$, the drift term $f$ is also learnable in DiffFlow, compared to the fixed liner function as in most diffusion models. Besides, these SDEs are jointly trained by minimizing the KL divergence. This allows the model to better adapt to changes in the data distribution, thus speeding up the convergence of the backward diffusion process. Similar to DiffFlow,  Zand et al. (2023) proposes a method called Diffusion with Normalizing Flow priors that also combines diffusion models with normalizing flows. The method first uses a linear SDE in the forward process to gradually convert the data distribution into a noise distribution. In the reverse process, a normalizing flow network is introduced to map the standard Gaussian distribution to latent variables close to the data distribution through a series of reversible transformations, which allows the samples to return to the data distribution more quickly, rather than relying on a large number of small incremental adjustments.
However, the fixed step sizes in existing SDE solvers Eq.(20), which usually require tremendous iterative steps, significantly affect generation efficiency. To address this,  Jolicoeur-Martineau et al. (2021) proposes a novel adaptive step-size SDE solver that dynamically adjusts the step size based on error tolerance, thereby reducing the number of evaluations. Specifically, the proposed method dynamically adjusts the step size by estimating the error between first-order and second-order approximations, leveraging a tolerance mechanism that incorporates both absolute and relative error thresholds. Furthermore, the use of extrapolation enhances precision without incurring additional computational overhead. This approach obviates the need for manual step-size tuning and is applicable across a range of diffusion processes, including Variance Exploding and Variance Preserving models. As a result of Gaussian assumption for reverse transition kernels becomes invalid when using limited sampling steps. The Gaussian Mixture Solver (GMS) (Guo et al., 2024) optimized SDE solver by using Gaussian mixture distribution. It addresses the limitations of the traditional process of SDE solvers in Eq.(21), which assume a Gaussian distribution for the reverse transition kernel. Specifically, GMS replaces the Gaussian assumption with a more flexible Gaussian mixture mode and utilizes a noise prediction network with multiple heads to estimate the higher-order moments of the reverse transition kernel. At each sampling step, it employs the Generalized Method of Moments to optimize the parameters of the Gaussian mixture transition kernel, allowing for a more accurate approximation of the true reverse process, even with a limited number of discretization steps.
Instead,  Xue et al. (2024a) unifies Bayesian Flow Networks (BFNs) with Diffusion Models (DMs) by introducing time-dependent SDEs into the BFN framework. BFNs work by iteratively refining the parameters of distributions at different noise levels through Bayesian inference, rather than directly refining the samples as in traditional diffusion models. To achieve theoretical unification between BFNs and DMs, the authors introduce a time-dependent linear SDE that governs the noise addition process in BFNs. This forward process includes two time-dependent functions: one controlling the drift of parameters and another controlling their diffusion. Based on this forward equation, they derive a corresponding reverse-time SDE for generating data from noise. This reverse process combines the drift term with a score-based correction term. This reverse-time SDE directly aligns with the denoising process in diffusion models, enabling the BFN sampling process to effectively replicate the behavior of diffusion models.
By optimizing the solving process of SDE in Eq.(20), Stochastic Adams Solver (SA-Solver) (Xue et al., 2024b) was presented. It is an innovative method designed to efficiently sample from Diffusion SDEs in Diffusion Probabilistic Models (DPMs) (Ho et al., 2020). By addressing the significant computational burden of traditional samplers, SA-Solver achieves this through a clever combination of variance-controlled diffusion SDEs and a stochastic Adams method (Buckwar & Winkler, 2006), which is a multi-step numerical technique that leverages prior evaluations to enhance efficiency. The method introduces a noise control function $\tau(t)$, enabling dynamic adjustment of the noise injected during sampling, which in turn strikes an optimal balance between sampling speed and the quality of the generated data. Operating within a predictor-corrector framework, SA-Solver first makes an initial estimate through the predictor step and then refines this estimate using the corrector step, ensuring greater accuracy in the final output. This strategic integration significantly reduces the number of function evaluations required.
ODE Solver. Unlike SDE solvers, the trajectories generated by ordinary differential equation (ODE) solvers are deterministic (Song et al., 2020c), remaining unaffected by stochastic variations. Consequently, these deterministic ODE solvers tend to achieve convergence more rapidly compared to their stochastic counterparts, although this often comes at the expense of a marginal reduction in sample quality. The corresponding deterministic process Eq.(24) can be derived from the reverse-time SDE Eq.(21) by removing the stochastic term $g(t)d\bar{\mathbf{w}}$, resulting in a deterministic process that shares the same marginal probability densities as the reverse-SDE:
$d\mathbf{x}=\left[{\mathbf{f}(\mathbf{x},t)}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\right]dt$ (24)
The forward process also exhibits a similar distinction between SDE and ODE approaches, yielding a deterministic process that preserves the same marginal distributions:
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t)dt$ (25)
Recent research has produced numerous works on faster diffusion samplers based on solving the ODE Eq.(24). Research shows that ODE samplers are highly effective when only a limited number of NFEs is available, while SDE samplers offer better resilience to prior mismatches (Nie et al., 2023) and exhibit superior performance with a greater number of NFEs (Lu et al., 2022).
Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a) builds upon the framework of Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), offering significant enhancements in sampling efficiency, which is one of the first models to leverage ODEs explicitly for the accelerating sampling process.
$q_{\sigma}(\mathbf{x}_{1:T}|\mathbf{x}_{0})=q_{\sigma}(\mathbf{x}_{T}|\mathbf{x}_{0})\prod_{t=2}^{T}q_{\sigma}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ (26)
Unlike DDPM’s Markovian forward process Eq.(1) where each state only depends on its immediate predecessor, DDIM utilizes the Non-Markovian Forward Process Eq.(26). These formulas allow each state not only to depend on its immediate predecessor but also on the initial state or a series of previous states. Specifically, Eq.(27) outlines how DDIM generates $\mathbf{x_{t-1}}$ from $\mathbf{x_{t}}$ by predicting the denoised observation, which essentially approximates reversing the diffusion process:
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})+\sigma_{t}\mathbf{\epsilon}_{t}$ (27)
During the process, DDIM employs an ODE solver to manage the continuous transformation across the latent space:
$d\mathbf{x}(t)=\epsilon_{\theta}^{(t)}\left(\frac{\mathbf{x}(t)}{\sqrt{\sigma^{2}+1}}\right)d\sigma(t)$ (28)
Eq.(28) is key to the efficient management of the generation process, allowing for fewer steps in the generative sequence by smoothly interpolating between states using an ODE solver, thus significantly reducing the time complexity compared to traditional methods.
While DDIM’s ODE formulation Eq.(24) and its implementation through Eq.(28) provide a foundation for deterministic sampling,  Liu et al. (2022a) identifies two critical issues in the ODE formulation of DDIM: first, the neural network $\theta$ and ODE are only well-defined within a narrow data manifold, while numerical methods generate samples outside this region. second, the ODE becomes unbounded as $t\to 0$ for linear schedules. Therefore PNDM is proposed to decompose the numerical solver into gradient and transfer components. It achieves second-order convergence, enabling 20x speedup while maintaining quality and reducing FID by  0.4 points at the same step count across different datasets and variance schedules.
The DPM-solver (Lu et al., 2022) and Diffusion Exponential Integrator Sampler (DEIS) (Zhang & Chen, 2022) innovate by leveraging the semi-linear structure of the probability flow ODE Eq.(24) to design custom ODE solvers that outperform traditional Runge-Kutta (Hochbruck & Ostermann, 2010) methods in terms of efficiency. Specifically, DPM-solver solves the linear part of the equation and uses neural networks to approximate the nonlinear component. Compared to PNDM, DPM-solver maintains lower FID scores at the same NFE. Further, DEIS employs an Exponential Integrator (Hochbruck & Ostermann, 2010) to discretize ODEs. This method simplifies the probability flow ODE by transforming the probability ODE into a simple non-stiff ODE. Both of the innovations reduce the number of iterations needed producing high-quality samples within just 10 to 20 iterations.
To reduce the computational overhead,  Zheng et al. (2023) presents an improved technique for maximum likelihood estimation of ODEs. Instead of directly working with the drift and score terms in Eq.(24), it introduces velocity parameterization to predict and optimize velocity changes $d\mathbf{x}_{t}$ during the diffusion process directly. The method improves upon previous ODE-based approaches by incorporating second-order flow matching for more precise trajectory estimation. Additionally, it introduces a negative log-signal-to-noise-ratio (log-SNR) for timing control of the diffusion process, alongside normalized velocity and importance sampling to reduce variance and optimize training. These enhancements significantly improve the model’s likelihood estimation performance on image datasets without variational dequantization or data augmentation. While previous methods focus on improving reverse ODE integrators based on Eq.(24), Denoising MCMC (DMCMC) (Kim & Ye, 2022) takes a different approach by integrating Markov Chain Monte Carlo (MCMC) with ODE integrators to optimize the data sampling process. In DMCMC, MCMC first generates initialization points in the product space of data and diffusion time, which are closer to a noise-free state, significantly reducing the noise levels that need to be processed by the ODE integrators. This hybrid approach complements rather than improves the ODE integrators directly, enhancing overall sampling efficiency.
Besides, Lu & Song (2024) focuses on improving continuous-time consistency models(CMs) (Song et al., 2023; Song & Dhariwal, 2023) for efficient diffusion sampling by modifying the ODE parameterization and training objectives of continuous-time CMs. The core contribution is TrigFlow, a unified framework that bridges EDM (Karras et al., 2022) and Flow Matching (Peluchetti, 2023; Lipman et al., 2022; Liu et al., 2022b; Albergo et al., 2023; Heitz et al., 2023).
While the traditional probability flow framework is governed by Eq.(24), they propose a simplified parameterization. To model these dynamics, they introduce a neural network $\mathbf{F_{\theta}}$ with parameters $\theta$ that takes normalized samples and time encodings as input. The time variable $t$ is transformed by $c_{noise}(t)$ to better condition the network. This results in a concise probability flow ODE:
$\frac{d\mathbf{x}_{t}}{dt}=\sigma_{d}\mathbf{F}_{\theta}(\frac{\mathbf{x}_{t}}{\sigma_{d}},c_{noise}(t))$ (29)
By introducing this simplified ODE parameterization, TrigFlow enables training large-scale CMs (up to 1.5B parameters) that achieve state-of-the-art performance with just two sampling steps, significantly reducing computational costs compared to DPM-solver (Lu et al., 2022) and other traditional diffusion models.
3.3.2 Sampling Scheduling
In diffusion models, a sampling schedule outlines a structured approach for timing and managing the sampling steps to improve both the efficiency and quality of the model’s output. It focuses on optimizing the sequence and timing of these steps, utilizing advanced techniques to process multiple steps simultaneously or in an improved sequential order. Specifically, this scheduling primarily targets the optimization of the reverse process in DDPM, as described in Eq.(4), where each step requires model prediction to gradually denoise from pure noise to the target sample. This scheduling is crucial for reducing computational demands and enhancing the model’s performance in generating high-quality samples.
Parallel Sampling. Parallel sampling is a process that schedules sampling tasks in parallel. Traditional diffusion models require a extensive series of sequential denoising steps to generate a single sample, which can be quite slow. For instance, Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) might need thousands of these steps to produce one sample. However, parallel sampling leverages the power of a multi-core GPU to compute multiple sampling steps. This approach optimizes the use of computational resources and reduces the time needed for model generation. Currently, there is significant work on autoregressive models that employ parallelization to speed up the sampling process. However, these techniques cannot be directly applied to diffusion models. This is because the computational frameworks and inference efficiency in autoregressive models differ from those in diffusion models. Therefore, designing algorithms tailored to parallelize the sampling process of diffusion models is crucial.
Figure 15: Computation graph of Picard iterations, which introduces skip dependencies (Shih et al., 2024).
An innovative extension of the Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020a) using Deep Equilibrium (DEQ) models is presented (Pokle et al., 2022), where the sampling sequence is conceptualized as a multivariate fixed-point system. This approach focuses on finding the system’s fixed point during the forward pass and utilizes implicit differentiation during the backward pass to enhance computational efficiency. By treating the sampling steps as an equilibrium system and solving for their fixed points simultaneously, parallel processing on multiple GPUs is achieved by batching the workload. Notably, it improves efficiency by updating each state $\mathbf{x}_{t}$ based on predictions from the noise prediction network $\epsilon_{\theta}$, which takes into account all subsequent states $\mathbf{x}_{t+1:T}$, unlike traditional diffusion processes that update states sequentially based only on the immediate next state $\mathbf{x}_{t+1}$.
ParaDiGMS (Shih et al., 2024) employs Picard iterations to parallelize the sampling process in diffusion models. This method models the denoising process using ordinary differential equations (ODEs) (Song et al., 2020c), where Picard iterations approximate the solution to these ODEs concurrently for multiple state updates. ParaDiGMS operates within a sliding window framework, enabling the simultaneous update of multiple state transitions. Each state is iteratively connected to different generations, allowing for information integration from several previous iterations Figure 15.
Building upon these parallel processing concepts, ParaTAA (Tang et al., 2024) also adopts an iterative approach, primarily applied in practical deployments for image generation tasks such as text-to-image transformations using Stable Diffusion. Specifically, ParaTAA enhances parallel sampling by solving triangular nonlinear equations through fixed-point iteration. Furthermore, the study introduces a novel variant of the Anderson Acceleration (Walker & Ni, 2011) technique, named Triangular Anderson Acceleration, designed to accelerate computation speed and improve the stability of iterative processes.
Timestep Schedule. In the sampling process of diffusion models, the entire process is discrete, and the model progressively restores data from noise through a series of discrete timesteps. Each timestep represents a small denoising step that moves the model from its current state closer to the real data. The timestep schedule refers to the strategy for selecting and arranging these timesteps.
Figure 16: Illustration of timestep schedule optimization process.
It may involve distributing them evenly or performing denser sampling during key stages to ensure the efficiency of the sampling process and the quality of the generated results. Selecting an appropriate method to choose a series of timesteps can enable the sampling process to converge quickly, which the process is shown in Figure 16.
FastDPM (Kong & Ping, 2021) is a unified framework for fast sampling in diffusion models that innovatively generalizes discrete diffusion steps to continuous ones and designs a bijective mapping between these continuous diffusion steps and noise levels. By utilizing this mapping, FastDPM constructs an approximate diffusion and reverse process, significantly reducing the number of steps required ($S\ll T$). It allows for the flexible determination of sampling points by selecting specific steps or variances from the original diffusion process, thereby enhancing efficiency. Watson et al. (2021) proposes a dynamic programming algorithm to optimize timestep scheduling in Denoising Diffusion Probabilistic Models (DDPMs). The algorithm efficiently determines the optimal timestep schedule from thousands of possible steps by leveraging the decomposable property of Evidence Lower Bound (ELBO) across consecutive timesteps and treating timestep selection as an optimization problem. Experiments show that the optimized schedule requires only 32 timesteps to achieve comparable performance to the original model with thousands of steps, effectively balancing efficiency and quality.
Figure 17: Minimizing an upper bound on the Kullback-Leibler divergence (KLUB) between the true and linearized generative SDEs to find optimal DM sampling schedules (Sabour et al., 2024).
However, optimizing an exact Evidence Lower Bound (ELBO) is typically not conducive to enhancing image quality. To address this, a universal framework named Align Your Steps (AYS) (Sabour et al., 2024)has been introduced, aimed at optimizing sampling schedules for various datasets, models, and stochastic SDE solvers. AYS uses stochastic calculus to optimize the sampling schedule allowing for the identification of optimal sampling timesteps by minimizing the Kullback-Leibler Divergence Upper Bound (KLUB) between discretized learnt SDE and true learnt SDE, which is shown in Figure 17.
3.3.3 Truncated Sampling
Truncated sampling enhances the efficiency of sample generation in diffusion models by strategically reducing redundant computations, thereby lowering computational costs while maintaining high-quality outputs. Methods like Early Exit focus on skipping unnecessary computations in later stages of the diffusion process when predictions are confident. Meanwhile, Retrieval-Guided Initialization improves efficiency in the early stages by leveraging retrieved examples to provide a better initialization, effectively skipping parts of the iterative refinement process. These approaches allocate computation more effectively by focusing resources on the most critical steps of the sampling process.
Early Exit. Early exit is a technique to improve the efficiency of large neural networks by reducing unnecessary computation. It allows intermediate layers to make early predictions, saving computational resources while maintaining performance. The core idea is to use intermediate predictions when they are confident enough, thereby skipping further layers for simpler inputs. Schuster et al. (2021) introduce Confident Adaptive Transformers (CATs). As shown in Figure 18, model $\mathcal{G}(X)$ provably consistent with the original $\mathcal{F}$ with arbitrarily high probability. They enhance transformer efficiency by adding prediction heads at intermediate layers and using a meta classifier to decide when to stop computation. This ensures high confidence in predictions, significantly improving efficiency on NLP tasks without sacrificing accuracy. Formally, they create a model $\mathcal{G}(X)$ that includes early classifiers $\mathcal{G}=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{l}\},\quad\text{ensuring}\quad P(\mathcal{G}(x)=\mathcal{F}(x))\geq 1-\epsilon$.
Figure 18: The CAT model $\mathcal{G}$ Schuster et al. (2021) improves computational efficiency by enabling early exits on certain inputs while ensuring predictive consistency with the full model $\mathcal{G}$.
Building on this, another work Moon et al. (2024) propose Adaptive Score Estimation (ASE) for diffusion models. ASE reduces sampling time by optimizing computation allocation at different time steps, introducing a time-dependent exit schedule. This method improves sampling efficiency while preserving image quality. The key idea is to dynamically allocate computation based on the difficulty of score estimation at each time step, allowing for early exits in simpler cases.
Retrieval-Guided Initialization. Retrieval-Guided Initialization combines the efficiency of retrieval mechanisms with the generative power of diffusion models. As illustrated in Figure 19, this approach begins with a retriever that selects relevant images from a database based on the input text prompts. These retrieved images, organized by specific keywords, serve as contextual guidance for the diffusion model, enhancing the relevance and quality of the generated output image.Blattmann et al. (2022) introduce a semi-parametric model that integrates neural network-based generative models with a retrieval mechanism, guided by examples from a large dataset. This approach shows the feasibility of combining retrieval with neural synthesis, although it requires a large dataset.  Sheynin et al. (2022) propose integrating a K-Nearest Neighbors (KNN) retrieval mechanism with diffusion models. By retrieving similar images from a large dataset, the model effectively guides the diffusion process, helping to generate high-quality images efficiently. This approach addresses the challenge of generating high-quality images with limited resources. Chen et al. (2022b) introduce a retrieval-augmented framework specifically for text-to-image generation, retrieving relevant images based on textual descriptions to guide the diffusion process and align the output with the input text.
Finally, Zhang et al. (2023b) introduces a learning-free inference mechanism for diffusion models. Instead of training complex networks, they retrieve and reuse diffusion trajectories from a precomputed database, significantly reducing the computational burden and providing an efficient solution for real-time applications.
Figure 19: Illustration of the retrieval-based diffusion model. The retriever selects relevant images from a database based on input text. These retrieved images provide contextual guidance for the generator (diffusion model) to produce a new, coherent output image.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=18em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ Compression, ver [ Quantization, fill=blue!10 [ Post-Training Quantization, fill=blue!10 [ PTQ4DM (Shang et al., 2023), Q-Diffsuion (Li et al., 2023), BRECQ (Li et al., 2021),
RAQ (Kim et al., 2024b), PTQD (He et al., 2024), leaf, text width=45em ] ] [ Quantization-Aware Training, fill=blue!10 [ TDQ (So et al., 2024), QALoRA (He et al., 2023), leaf, text width=45em ] ] ] [ Pruning, fill=blue!10 [ Diff-Pruning (Fang et al., 2023), LD-Pruner (Castells et al., 2024), LayerMerge (Kim et al., 2024a), LAPTOPDiff (Zhang et al., 2024), leaf, text width=61.7em ] ] [ Knowledge Distillation, fill=blue!10 [ Vector Field Distillation, fill=blue!10 [ Denoising Student (Luhman & Luhman, 2021), Progressive Distillation (Salimans & Ho, 2022),
 Meng et al. (2023), CM (Song et al., 2023), leaf, text width=45em ] ] [ Generator Distillation, fill=blue!10 [ SDS (Poole et al., 2022), VSD (Wang et al., 2024b), Diff-Instruct (Luo et al., 2024),
CSD (Decatur et al., 2024), leaf, text width=45em ] ] ] ]
Figure 20: Summary of compression techniques for DMs.
3.4 Compression
Model compression enhances efficiency by reducing the sizes and the amount of arithmetic operations of DM. As summarized in Figure 28, model compression techniques for DMs can be grouped into three categories: quantization, pruning, and knowledge distillation. These three categories are orthogonal to each other, and compress DMs from different perspectives.
3.4.1 Quantization
Quantization compresses neural networks by converting model weights and/or activations of high-precision data types $\mathbf{X}^{\mathrm{H}}$ such as 32-bit floating point into low-precision data types $\mathbf{X}^{\mathrm{L}}$ such as 8-bit integer (Dettmers et al., 2024). Quantization techniques can be classified into post-training quantization (PTQ) and quantization-aware training (QAT).
Figure 21: Illustrations of the quantization.
Post-Training Quantization. PTQ involves selecting operations for quantization, collecting calibration samples, and determining quantization parameters for weights and activations. While collecting calibration samples is straightforward for CNNs and ViTs using real training data, it poses a challenge for Diffusion Models (DMs). In DMs, the inputs are generated samples $\mathbf{x}_{t}$ at various time steps (t = 0, 1, …, T), where T is large to ensure convergence to an isotropic Normal distribution. To address this issue, Shang et al. (2023) proposes PTQ4DM, the first DM-specific calibration set collection method, generating calibration data across all time steps with a specific distribution. However, their explorations remain confined to lower resolutions and 8-bit precision. Q-Diffsuion (Li et al., 2023) propose a time step-aware calibration data sampling to improve calibration quality and apply BRECQ (Li et al., 2021), which is a commonly utilized PTQ framework, to improve performance. Furthermore, compared to conventional PTQ calibration methods, they identify the accumulation of quantization error across time steps as another challenge in quantizing DMs Figure 22 (a). Therefore, they also propose a specialized quantizer for the noise estimation network shown in Figure 22 (b). Based on Q-Diffusion, Kim et al. (2024b) find that inaccurate computation during the early stage of the reverse diffusion process has minimal impact on the quality of generated images. Therefore, they introduce a method that focuses on further reducing the number of activation bits for the early reverse diffusion process while maintaining high-bit activations for the later stages. Lastly, He et al. (2024) presents PTQD, a unified formulation for quantization noise and diffusion perturbed noise. Additionally, they introduce a step-aware mixed precision scheme, which dynamically selects the appropriate bitwidths for synonymous steps.
Figure 22: Traditional PTQ scenarios and Q-Diffusion differ in (a) the creation of calibration datasets and (b) the workflow for model inference  (Li et al., 2023).
Quantization-Aware Training. Different from PTQ, QAT quantizes diffusion models during the training process, allowing models to learn quantization-friendly representations. Since QAT requires additional training after introducing quantization operators, it is much more expensive and time-consuming than PTQ. So et al. (2024) proposes a novel quantization method that enhances output quality by dynamically adjusting the quantization interval based on time step information. The proposed approach integrates with the Learned Step Size Quantization (Esser et al., 2019) framework, replacing the static quantization interval with a dynamically generated output from the Time-Dynamic Quantization module. This dynamic adjustment leads to significant improvements in the quality of the quantized outputs. He et al. (2023) introduces a quantization-aware low-rank adapter that integrates with model weights and is jointly quantized to a low bit-width. This approach distills the denoising capabilities of full-precision models into their quantized versions, utilizing only a few trainable quantization scales per layer and eliminating the need for training data.
3.4.2 Pruning
Pruning compresses DMs by removing redundant or less important model weights. Currently, most pruning methods for DMs focus on pruning structured patterns such as groups of consecutive parameters or hierarchical structures. For instance, Diff-Pruning (Fang et al., 2023) introduces the first dedicated method designed for pruning diffusion models. Diff-Pruning leverages Taylor expansion over pruned timesteps to estimate the importance of weights. By filtering out non-contributory diffusion steps and aggregating informative gradients, Diff-Pruning enhances model efficiency while preserving essential features.
LD-Pruner (Castells et al., 2024), as illustrated in Figure 23, on the other hand, proposes a pruning method specifically designed for Latent Diffusion Models (LDMs) The key innovation of LD-Pruner lies in its utilization of the latent space to guide the pruning process. The method enables a precise assessment of pruning impacts by generating multiple sets of latent vectors—one set for the original Unet and additional sets for each modified Unet where a single operator is altered. The importance of each operator is then quantified using a specialized formula that considers shifts in both the central tendency and variability of the latent vectors. This approach ensures that the pruning process preserves model performance while adapting to the specific characteristics of LDMs.
Figure 23: Pruning evaluates changes in the central tendency and variability to determine the significance of each operator.  (Castells et al., 2024).
Kim et al. (2024a) introduces a technique known as LayerMerge, designed to jointly prune convolution layers and activation functions to achieve a desired inference speedup while minimizing performance degradation. LayerMerge addresses the challenge of selecting which layers to remove by formulating a new surrogate optimization problem. Given the exponential nature of the selection space, the authors propose an efficient solution using dynamic programming. Their approach involves constructing dynamic programming (DP) lookup tables that exploit the problem’s inherent structure, thereby allowing for an exact and efficient solution to the pruning problem.
Figure 24: Illustrations of the pruning.
Lastly, LAPTOPDiff (Zhang et al., 2024) introduces a layer-pruning technique aimed at automatically compressing the U-Net architecture of diffusion models. The core of this approach is an effective one-shot pruning criterion, distinguished by its favorable additivity property. This property ensures that the one-shot performance of the pruning is superior to other traditional layer pruning methods and manual layer removal techniques. By framing the pruning problem within the context of combinatorial optimization, LAPTOPDiff simplifies the pruning process while achieving significant performance gains. The proposed method stands out for its ability to provide a robust one-shot pruning solution, offering a clear advantage in compressing diffusion models efficiently.
3.4.3 Knowledge Distillation
Knowledge distillation (Hinton et al., 2015) is a technique that compresses complex models into smaller, efficient versions with minimal performance loss. The process of knowledge distillation can be captured by minimizing the following loss function:
$L_{\text{KD}}=\alpha L_{\text{CE}}(y,\sigma(T_{s}(x)))+\beta L_{\text{MSE}}(T_{t}(x),T_{s}(x)),$ (30)
Figure 25: Illustrations of the knowledge distillation.
where $T_{t}$ and $T_{s}$ are the teacher and student models, respectively, $\sigma$ is the softmax function, $L_{\text{CE}}$ is the cross-entropy loss, and $L_{\text{MSE}}$ is the mean squared error loss, with $\alpha$ and $\beta$ as balancing hyperparameters. In DMs, known for generating high-quality data, this approach is increasingly applied to improve efficiency by addressing slow sampling speeds caused by the numerous neural function evaluations in the diffusion process. By distilling the knowledge from DMs into more efficient forms, researchers aim to accelerate sampling while preserving the generative performance of the original models. Follow Luo (2023), knowledge distillation for DMs can be categorized into vector field distillation and generator distillation.
Vector Field Distillation. Vector field distillation improves the efficiency of deterministic sampling in diffusion models by transforming the generative ODE into a new generative vector field. This approach reduces the number of NFEs needed to produce samples of similar quality. Luhman & Luhman (2021) first proposes a strategy to distill a DDIM sampler into a Gaussian model that needs only one NFE for sampling. In this approach, a conditional Gaussian model serves as the student model, and the training process involves minimizing the conditional KL divergence between this student model and the DDIM sampler.
Figure 26: The progressive distillation, where the original sampler derived from integrating a learned diffusion model’s probability flow ODE, is efficiently condensed into a new sampler that achieves the same task in fewer steps.  (Salimans & Ho, 2022).
While this method advances the application of knowledge distillation to diffusion models, it still has computational inefficiencies, as it necessitates generating the final outputs of DDIM or other ODE samplers, which entails hundreds of NFEs for each training batch. Salimans & Ho (2022) proposes a progressive distillation strategy to train a student model to use half the NFEs of the teacher model by learning its two-step prediction strategy, as illustrated in Figure 26. Once the student model accurately predicts the teacher’s two-step sampling strategy, it replaces the teacher model, and a new student model is trained to further reduce the sampling steps. This method reduces the NFEs significantly, achieving 250 times greater efficiency with only a 5% drop in performance.
A two-stage distillation strategy is proposed by Meng et al. (2023) to address the challenge of transferring knowledge from classifier-free guided conditional diffusion models like DALL·E-2 (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022). In the first stage, a student model is trained with classifier-free guidance to learn from the teacher diffusion model. The second stage employs the progressive diffusion strategy to further reduce the number of diffusion steps for the student model. This two-stage approach is applied to both pixel-space and latent-space models for various tasks, including text-guided generation and image inpainting.  Song et al. (2023) introduce the Consistency Model (CM), which leverages the self-consistency property of generative ODEs in diffusion models. Instead of directly mimicking the output of the generative ODE, their method focuses on minimizing the difference in the self-consistency function. By randomly diffusing a real data sample and simulating a few steps of the generative ODE to generate another noisy sample on the same ODE path, the model inputs these two noisy samples into a student model.
Generator Distillation. Unlike vector field distillation, which primarily focuses on distilling knowledge into student models with identical input and output dimensions, generator distillation aims to transfer the complex distributional knowledge embedded in a diffusion model into a more efficient generator. The Neural Radiance Field (NeRF) (Mildenhall et al., 2021) is a powerful technique for reconstructing 3D scenes from 2D images by learning a continuous volumetric scene representation. NeRFs generate photorealistic views of scenes from novel angles, making them valuable for applications in computer vision and graphics.
Figure 27: Illustration as  (Poole et al., 2022), it utilizes score distillation sampling.
However, the limited availability of data for constructing NeRFs is an issue. Therefore, exploring distillation methods to obtain NeRFs with contents related to given text prompts is a promising way. (Poole et al., 2022) first proposed Score Distillation Sampling (SDS) to distill a 2D text-to-image diffusion model into 3D NeRFs, as illustrated in Figure 27. Unlike traditional NeRF construction that requires images from multiple views of the target 3D objects, text-driven construction of NeRF lacks both the 3D object and the multiple views. The SDS method optimizes the NeRF by minimizing the diffusion model’s loss function using NeRF-rendered images from a fixed view.
Wang et al. (2024b) introduce Variational Score Distillation (VSD), which extends SDS by treating the 3D scene corresponding to a textual prompt as a distribution rather than a single point. Compared to SDS, which generates a single 3D scene and often suffers from limited diversity and fidelity, VSD is capable of generating more varied and realistic 3D scenes, even with a single particle.  Luo et al. (2024) propose Diff-Instruct, which can transfer knowledge from pre-trained diffusion models to a wide range of generative models, all without requiring additional data. The key innovation in Diff-Instruct is the introduction of Integral Kullback-Leibler divergence, which is specifically designed to handle the diffusion process and offers a more robust way to compare distributions.  Decatur et al. (2024) present Cascaded Score Distillation (CSD), an advancement by addressing a key limitation of standard SDS. Specifically, while traditional SDS only leverages the initial low-resolution stage of a cascaded model, CSD distills scores across multiple resolutions in a cascaded manner, allowing for nuanced control over both fine details and the global structure of the supervision. By formulating a distillation loss that integrates all cascaded stages, which are trained independently, CSD enhances the overall capability of generating high-quality 3D representations.
4 System-Level Efficiency Optimization
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=22em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ System, ver [ Optimized Hardware-Software Co-design, fill=blue!10 [  Chen et al. (2023c), SDA (Yang et al., 2023a),  Choi et al. (2024a), leaf, text width=61.7em ] ] [ Parallel Computing, fill=blue!10 [ DistriFusion (Li et al., 2024a), PipeFusion (Wang et al., 2024a), SwiftDiffusion (Li et al., 2024b), DiffusionPipe (Tian et al., 2024), leaf, text width=61.7em ] ] [ Caching Technique, fill=blue!10 [ NIRVANA (Agarwal et al., 2024), DeepCache (Ma et al., 2024b), Block Caching (Wimbauer et al., 2024), L2C (Ma et al., 2024a),
FORA (Selvaraju et al., 2024), leaf, text width=61.7em ] ] ]
Figure 28: Summary of system-level efficiency optimization techniques for diffusion models.
4.1 Hardware-Software Co-Design
The co-design of hardware and software is pivotal for achieving efficient deployment of diffusion models in real-time and resource-constrained environments. Following algorithm-level optimizations, system-level techniques focus on integrating hardware-specific features, distributed computation, and caching mechanisms. These strategies aim to address the computational complexity and memory demands of large-scale diffusion models, enabling more practical applications across various platforms like GPUs, FPGAs, and mobile devices. One significant contribution is the work by Chen et al. (2023c), which explores GPU-aware optimizations for accelerating diffusion models directly on mobile devices. Implementing specialized kernels and optimized softmax operations reduces inference latency, achieving near real-time performance on mobile GPUs. In a related effort,  Yang et al. (2023a) propose SDA, a low-bit stable diffusion accelerator designed specifically for edge FPGAs. Utilizing quantization-aware training and a hybrid systolic array architecture as illustrated in Figure 29, SDA effectively balances computational efficiency with flexibility, handling both convolutional and attention operations efficiently. Through a two-level pipelining structure, the nonlinear operators are efficiently integrated with the hybridSA, enabling coordinated operation that enhances processing speed while reducing resource usage. Finally, SDA achieves a speedup of 97.3x when compared to ARM Cortex-A53 CPU.
Furthermore, Choi et al. (2024a) introduces a stable diffusion processor optimized for mobile platforms through patch similarity-based sparsity, mixed-precision strategies and and a Dual-mode Bit-Slice Core (DBSC) architecture that supports mixed-precision computation, which particularly targeting resource-constrained devices such as mobile platforms. Together, these optimizations significantly improve throughput and energy efficiency, making Stable Diffusion more viable for energy-sensitive applications.
4.2 Parallel Computing
Figure 29: Illustration of the HybridSA architecture from  Yang et al. (2023a).
Parallel computing plays a critical role in the efficient execution of diffusion models, especially given the computation-intensive nature of these algorithms in Figure 30. Recent advances in parallel computing strategies have enabled significant improvements in inference speed and scalability, often without compromising the quality of the generated output (Li et al., 2024a; Wang et al., 2024a; Li et al., 2024b; Tian et al., 2024). This section highlights several notable contributions that tackle the challenge of parallelizing diffusion models across multiple GPUs and other distributed architectures. Li et al. (2024a) introduced DistriFusion, a framework designed for distributed parallel inference tailored to high-resolution diffusion models such as SDXL. Their approach involves partitioning the model inputs into distinct patches, which are then processed independently across multiple GPUs. This method leverages the available hardware resources more effectively, achieving a 6.1x speedup on 8xA100 GPUs compared to single-card operation, all while maintaining output quality.
To address potential issues arising from the loss of inter-patch interaction, which could compromise global consistency, DistriFusion employs dynamic synchronization of activation displacements, striking a balance between preserving coherence and minimizing communication overhead.
Figure 30: Illustrations of the parallel computing for diffusion models.
Building on the insights gained from DistriFusion, Wang et al. (2024a) further refined the distributed inference paradigm with PipeFusion. This system not only splits images into patches but also distributes the network layers across different devices, thereby reducing the associated communication costs and enabling the use of PCIe-linked GPUs instead of NVLink-connected ones. PipeFusion integrates sequence parallelism, tensor parallelism, displaced patch parallelism, and displaced patch pipeline parallelism, optimizing workflow for a wider range of hardware configurations.
Figure 31: Illustrations of the diffusion architecture from (Li et al., 2024a).
For applications involving add-on modules such as ControlNet and LoRA, Li et al. (2024b) developed SwiftDiffusion, as illustrated in Figure 31. This framework optimizes the serving workflow of these modules, allowing them to run in parallel on multiple GPUs. As a result, SwiftDiffusion delivers a 5x reduction in inference latency and a 2x improvement in throughput, ensuring that enhanced speed does not come at the expense of output quality.
Lastly, Tian et al. (2024) focused on the training phase with DiffusionPipe, demonstrating that pipeline parallelism can produce a 1.41x training speedup, while data parallelism contributes an additional 1.28x acceleration. Although the optimization methods for DiffusionPipe were not detailed in the notes, the combination of these parallelization strategies offers a promising direction to improve the efficiency of both the training and inference pipelines for diffusion models.
4.3 Caching Technique
In diffusion models, the computational hotspot often centers around discrete time-step diffusion, which is characterized by strong temporal locality. Consequently, building an efficient caching system for diffusion models is nonnegligible to enhance its performance. Indeed, extensive research has been conducted on optimizing caching systems in Figure 32, resulting in significant advancements in this field.
Agarwal et al. (2024) proposed NIRVANA, a novel system designed to enhance the efficiency of text-to-image generation using diffusion models. Specifically, the key innovation lies in its approximate caching technique, which reduces computational costs and latency by reusing intermediate noise states from previous image generation processes. Instead of starting from scratch with every new text prompt, NIRVANA retrieves and reconditions these cached states, allowing it to skip several initial denoising steps.
Figure 32: Illustrations of the caching system for diffusion models focus on the U-Net block and the Transformer layer, critical components for effectively implementing caching techniques.
Additionally, the system uses a custom cache management policy called Least Computationally Beneficial and Frequently Used (LCBFU), which optimizes the storage and reuse of cached states to maximize computational efficiency. This makes NIRVANA particularly suited for large-scale, production-level deployments of text-to-image diffusion models.
From another perspective, Ma et al. (2024b) introduces an innovative approach called DeepCache, designed to accelerate the image generation process by leveraging the temporal redundancy in the denoising steps of diffusion models, without the need for additional model training, as illustrated in Figure 33. The key insight is the observation that high-level features, such as the main structure and shape of an image, exhibit minimal changes between adjacent denoising steps. These features can be cached and reused in subsequent steps, thereby avoiding redundant computations. This method takes advantage of the U-Net architecture by combining these cached high-level features with low-level features, updating only the low-level features to reduce computational load, leading to a significant acceleration in the overall process.  Wimbauer et al. (2024) proposed Block Caching, a technique that identifies and caches redundant computations within the model’s layers during the denoising process. By reusing these cached outputs in subsequent timesteps, the method significantly speeds up inference while maintaining image quality. To optimize this caching process, they introduce an Automatic Cache Scheduling mechanism, which dynamically determines when and where to cache based on the relative changes in layer outputs over time. Additionally, the paper addresses potential misalignment issues from aggressive caching by implementing a Scale-Shift Adjustment mechanism, which fine-tunes cached outputs to align with the model’s expectations, thereby preventing visual artifacts.
Figure 33: Illustration of the caching system from (Ma et al., 2024b).
Recently, the application of diffusion with transformer models has yielded considerable success. Ma et al. (2024a) is concerned with the introduction of a layer caching mechanism, designated Learning-to-Cache (L2C), to accelerate diffusion transformer models. L2C exploits the redundancy between layers within the transformer architecture, dynamically caching computations from certain layers to reduce redundant calculations and lower inference costs. The implementation entails transforming the layer selection problem into a differentiable optimization problem, using interpolation to determine whether to perform a full computation or utilize cached results at different timesteps during inference. In contrast to the emphasis on layer caching,  Selvaraju et al. (2024) proposed Fast-Forward Caching (FORA), a technique designed to accelerate Diffusion Transformers (DiT) by reducing redundant computations during the inference phase. The key insight behind FORA is the observation that the outputs from the self-attention and MLP layers in a Transformer exhibit high similarity across consecutive time steps in the diffusion process. To leverage this, FORA implements a static caching mechanism where these layer outputs are cached at regular intervals, which are determined by $N$, and reused for a set number of subsequent steps, thereby avoiding recomputing similar outputs.
5 Frameworks
Diffusion model frameworks can generally be categorized based on their support for tasks such as training, fine-tuning, and inference. Specifically, frameworks that support training and/or fine-tuning are designed to provide scalable, efficient, and flexible infrastructures that enhance computational efficiency, reduce memory usage, and ensure the reliability of the training and fine-tuning processes. On the other hand, frameworks focused on inference aim to optimize throughput and reduce latency, addressing the practical needs of real-world applications. These frameworks offer various deployment options to handle diverse diffusion model tasks. Table 2 provides a summary of existing diffusion model frameworks along with their key features.
Table 2: Comparison of diffusion model frameworks.
Framework Training Fine-Tuning Inference Key Features
Diffusers Multi-model, versatile, open-source.
DALL-E × × Closed-source, API, high-quality.
OneDiff × × Optimized loading, CLI support.
LiteGen Efficient training, multi-GPU.
InvokeAI × Stable Diffusion, CLI, API.
ComfyUI-Docker × Modular, multiple models.
Grate × Image grids, model merging.
Versatile Diffusion Multi-modal, modular design.
UniDiffuser Unified, multi-modal diffusion.
Diffusers is an open-source framework by Hugging Face that provides a versatile toolset for working with diffusion models. It supports multi-model usage for text-to-image and image editing tasks. By leveraging a variety of pre-trained models and pipelines, it enables quick experimentation and fine-tuning for both research and production environments. The framework is designed to be extensible, supporting efficient model training, fine-tuning, and inference.
DALL-E is a closed-source text-to-image model developed by OpenAI. The framework is accessible via the OpenAI API, enabling users to generate high-quality images based on textual descriptions. Although it does not support direct training or fine-tuning, it provides inference capabilities that deliver state-of-the-art image synthesis. Its commercial use is constrained by the terms of the OpenAI API.
OneDiff focuses on optimizing model loading and inference processes, particularly in environments with limited computational resources. It is designed to offer fast execution and support for command-line interface (CLI) operations. OneDiff provides a way to quickly perform inference using various diffusion models, making it suitable for both developers and researchers who need a streamlined workflow.
LiteGen is a lightweight and efficient training framework tailored for diffusion models. It offers multi-GPU support and optimizes various tasks, enabling users to fine-tune pre-trained models with low computational overhead. LiteGen is especially useful for multi-task scenarios, where training speed and resource efficiency are critical.
InvokeAI is designed to facilitate both fine-tuning and inference processes. It includes support for multiple diffusion models and provides both a command-line interface (CLI) and an API for ease of use. InvokeAI is aimed at users who need a flexible and accessible toolkit for experimenting with image generation.
ComfyUI-Docker is a modular framework that supports multiple diffusion models like Stable Diffusion and ControlNet. While it offers a graphical user interface through Docker, it also allows for command-line operations. The modular design enables users to integrate various models and pipelines, providing a versatile platform for different image generation tasks.
Grate is a diffusion model toolkit specialized in generating image grids using multiple Stable Diffusion models. It supports model merging and fine-tuning to achieve varied artistic effects. Geared towards both artists and researchers, Grate offers flexibility in combining models to explore diverse image generation possibilities.
Versatile Diffusion is a framework that emphasizes multi-modal tasks and modular design. It supports training, fine-tuning, and inference, allowing users to build customized workflows tailored to specific domains. Its extensibility makes it suitable for a wide range of applications, from text-to-image generation to image-to-image translation.
UniDiffuser provides a unified framework for multi-modal diffusion, enabling text-to-image, image-to-text, and other tasks in a seamless manner. It supports training, fine-tuning, and inference, with a focus on bridging the gap between different types of data modalities. UniDiffuser is ideal for researchers who need a flexible model to handle complex diffusion tasks across various input forms.
6 Conclusion
In this survey, we provide a systematic review of efficient diffusion models, an important area of research aimed at democratizing diffusion models. We start with motivating the necessity for efficient diffusion models. Guided by a taxonomy, we review efficient techniques for diffusion models from algorithm-level and system-level perspectives respectively. Furthermore, we review diffusion models frameworks with specific optimizations and features crucial for efficient diffusion models. We believe that efficiency will play an increasingly important role in diffusion models and diffusion models-oriented systems. We hope this survey could enable researchers and practitioners to quickly get started in this field and act as a catalyst to inspire new research on efficient diffusion models.